. A 980 kg car traveling at 20 m/s rounds a curve of radius 40 m. What is the friction force that must act on the car to keep it in its circular path?
1
Expert's answer
2012-11-19T08:19:08-0500
Friction force is equal to centrifugal in this case
F_f=F_c = mv^2/r = 980*20^2/40 =9800 N